How do I solve -4xy where x=1/2 and y=-3?
Answer by Alan3354(69443)

How do I solve -4xy where x=1/2 and y=-3?
------------
You don't.
Equations are solved.
You can evaluate, or find out what -4xy is when x=1/2 and y=-3.
Sub 1/2 for x and =3 for y.